OpenStudy (nuccioreggie):

A pair of equations is shown below. x + y = 5 y = one halfx + 2 If the two equations are graphed, at what point do the lines representing the two equations intersect? (2, 5) (5, 2) (2, 3) (3, 2)

pooja195 (pooja195):

Graph the two problems here: https://www.desmos.com/calculator And figure out where they intersect, you can also solve by using substitution

OpenStudy (michele_laino):

according to the observation of @pooja195 if I substitute the value of \(y\) from second equation, into the first one, I get: \[\huge x + \left( {\frac{x}{2} + 2} \right) = 5\] please solve for \(x\)

OpenStudy (nuccioreggie):

x=2

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

now, please replace \(x=2\) into the second equation, you should get this: \[\huge y = \frac{x}{2} + 2 = \frac{2}{2} + 2 = ...?\]

OpenStudy (nuccioreggie):

b

OpenStudy (michele_laino):

in option B, we have \(x=5\), and not \(x=2\)

OpenStudy (nuccioreggie):

so whats the correct answer

OpenStudy (michele_laino):

no, please here is another hint: \[\huge y = \frac{x}{2} + 2 = \frac{2}{2} + 2 = 1 + 2 = ...?\]

OpenStudy (michele_laino):

so, \(y=...?\)

OpenStudy (nuccioreggie):

A

OpenStudy (michele_laino):

I got \(y=3\), so, what is the right option?

OpenStudy (michele_laino):

if \(x=2\) and \(y=3\) then the solution is given by the ordered pair \((2,3)\)

OpenStudy (nuccioreggie):

c

OpenStudy (michele_laino):

that's right!